# BitDecoding: Unlocking Tensor Cores for Long-Context LLMs with Low-Bit KV Cache

### v2

@@ IV-A Methods for optimizing low-bit layout on Tensor Cores
(1) Inducing low-bit optimized layout with hardware instructions. Our design is motivated by a novel insight: the thread-to-register mapping of ldmatrix loads data in Tensor Core’s interleaved fragment layout. As shown in Fig. 5-(2), if each thread then quantizes and packs locally, the resulting low-bit packing implicitly preserves the half-precision (FP16) interleaved layout. On unpacking and dequantization, values already match Tensor Core registers—no global reshape is required. Thus, rather than relying on heavyweight global transforms via manual implementations [8] or iterative search [31] as in prior methods, we use hardware instructions to automatically induce a valid low-bit packing layout while computing. This yields zero-overhead remapping that is efficient, compatible with Tensor Cores execution, and avoids extra data movement.
@@
To mitigate this inefficiency, we further design a faster dequantization mapping approach based on low-level bitwise operations and instructions inspired by [13]. After loading packed data into registers using ldmatrix, we cast them to INT32 before mapping them to the interleaved Tensor Core layout following the 75316420 pattern. This layout enables efficient conversion of INT4/INT2 data to FP16 using the lop3 instruction for bitwise manipulation while aligning with the Tensor Core computation pattern.
@@
(5) Unlocking warpgroup acceleration capabilities via smart uses of PTX-level instructions. GPUs, such as Hopper Tensor Cores, increasingly introduce Warpgroup Matrix Multiply-Accumulate (wgmma) instruction. This instruction however imposes a key constraint: in a matrix multiplication $C=AB$, only $A$ and $C$ can be sourced from registers, while $B$ must reside in shared memory. This presents a challenge for low-bit quantized data, as values are typically upconverted to FP16 in registers before computation. To resolve this, we leverage Hopper’s STSM PTX instruction to store dequantized FP16 values in shared memory efficiently, accessible for wgmma_SS operations. Remarkably, the asynchronous nature of WGMMA overlaps storage with computation, optimizing performance.
IV-B Strategies for parallelizing warps in dequantization
The second challenge is ensuring BitDecoding avoids the pitfalls of existing warp-parallelization strategies for mixed-precision attention, which suffer from low hardware utilization due to frequent warp stalls. Our key insight is that low-bit data moves at much higher bandwidth than full precision, shifting the bottleneck from memory to compute. We therefore design a warp layout that exploits the GPU memory hierarchy to parallelize dequantization efficiently, minimizing data movement and substantially improving Tensor Cores utilization (Table III demonstrates minimal overhead).
(1) Enhancing warps paralleism for dequantization. We introduce a novel warps layout to enable parallel dequantization of multiple packed data chunks. As illustrated in Fig. 6, unlike the original warp partitioning strategy, which allocates multiple warps along the $M$ dimension, we first reduce the warp allocation along $M$ to $W_{m}=1$—since the query length is typically less than 16—while increasing the number of warps along the $N$ dimension ($W_{n}$).
By increasing $W_{n}$, dequantization stalls can be effectively mitigated by the Streaming Multiprocessor (SM) warp scheduler [22], as multiple warps concurrently execute dequantization on packed data before proceeding to Tensor Cores-based matrix multiplication.
@@ V System Implementation
In this section, we describe how we implement BitDecoding, as illustrated in Fig. 7. Our implementation consists of three major components: (i) a query transformation component that supports diverse attention variants in LLMs; (ii) a Residual Kernel that performs low-cost quantization and packing while remaining general to both tensor-wise and channel-wise scaling across quantization algorithms; and (iii) a Packing Kernel with a fine-grained pipeline that fully utilizes both Tensor Cores and CUDA cores.
@@ V-A Query Transformation
Modern LLMs adopt diverse attention variants [9, 32, 16] with different key–value (KV) sharing patterns. BitDecoding aims to support all these variants.
@@ V-C Packing Kernel
(1) Optimizing asynchronous data movement. From Global to Shared Memory, we follow FlashAttention [5] via block-wise tiling [30] and strategic recomputation. It processes input matrices $Q\in\mathbb{R}^{T_{m}\times d}$, $K,V\in\mathbb{R}^{T_{n}\times d}$ in tiles within shared memory, using block sizes $T_{m}$ and $T_{n}$. The number of key-value tiles is $C_{n}=\lceil L/T_{n}\rceil$.
@@
From Shared Memory to Register, we use the PTX instruction ldmatrix to efficiently load $K_{\text{pack}}$, $V_{\text{pack}}$ and $sAcc$ from shared memory into registers with the Tensor Cores tiling layout. To eliminate bank conflicts, we use a sizzling scheme [4] defined as:
@@
Figure 8: Kernel performance on RTX4090.
@@ VI Evaluation
In this section, we comprehensively evaluate BitDecoding against state-of-the-art approaches and systems. Our evaluation has three primary goals:
@@
    Micro-benchmarking the attention kernels using low-bit KV cache on different GPU architectures (Nvidia 4090, A100 and H100) (Section VI-A).
@@
    Evaluating the end-to-end LLMs inference performance with low-bit KV cache (Section VI-B).
@@
    Analyzing the accuracy impact of low-bit KV cache, performance overhead of design, along with a breakdown of the performance contributions from individual components within BitDecoding (Section VI-C).
Figure 9: Kernel performance on A100.
@@ VI-A Kernels Performance Across GPU Architectures
  • Page: A high-throughput scenario where a larger batch_size is managed using the page management technique [14].
Baselines. We compare BitDecoding against several representative attention kernel implementations. For FP16 KV cache, we use FlashDecoding [5, 23]—a split-partitioned variant of FlashAttention optimized for long-context decoding—as our baseline for speedup normalization. For low-bit KV cache, we evaluate Kivi [17], a non-fused kernel supporting 4-bit and 2-bit quantization; Atom [35] and QServe [15], both fused-kernel implementations with CUDA Cores-only approach and supporting 4-bit cache with page management. Notably, Atom does not support GQA.
@@
Results on Bandwidth-constrained GPU (RTX 4090). Leveraging low-precision data types is a key strategy for accelerating data movement, particularly on GPUs with limited memory bandwidth. As shown in Fig. 8, under Single and Batches serving settings, BitDecoding-4bit achieves roughly $4\times$ speedup over FlashDecoding-v2, with the 2-bit implementation exceeding $7\times$. This significant improvement comes from its low-bit KV Cache, alleviating memory bottlenecks due to limited DRAM bandwidth on the RTX 4090.
For low-bit KV cache, BitDecoding significantly outperforms both non-fused and fused baselines across all scenarios. Compared to the non-fused KIVI, which is implemented using separate kernels, BitDecoding achieves dramatically higher speedups due to its fully fused design and optimized memory access patterns. In contrast, KIVI exhibits limited speedup in MHA and severe degradation in GQA.
In the Page setting, BitDecoding also outperforms fused implementations. In the MHA case (top right), BitDecoding-KC-4 achieves over $6\times$ speedup, while QServe and Atom reach only 3.5× and 2.5×, respectively. The advantage is even more pronounced in GQA (bottom right), where BitDecoding-KC-4 and KC-2 reach up to $3\times$ speedup, while QServe drops to just 1.4×. This highlights BitDecoding’s effective exploitation of Tensor Cores, enabling consistent acceleration across attention variants. In particular, for compute-intensive patterns like GQA, BitDecoding maintains high performance, whereas CUDA Cores-only baselines suffer significant degradation.
@@ VI-C Accuracy, Overhead and Performance Breakdown
TABLE I: Efficiency and accuracy tradeoff with low-bit KV cache. We use Llama-3.1-8B-Instruct with seq_len=32Kseq\_len=32K, and evaluate average accuracy on longbench [2].
@@
Figure 13: Runtime overhead of the residual KV cache.
Quantization and Packing Overhead. We evaluate the latency of quantization and packing under a sequence length of seq_len=128Kseq\_len=128K, comparing BitDecoding with Marlin [8] and Ladder [31]. As shown in Table LABEL:tab:quantization, the pre-transformation and packing step in previous mixed-precision computing methods introduce significant overhead, which cannot be ignored. Our kernel incurs minimal overhead after the Prefill phase, primarily due to kernel launch overhead. Moreover, during decoding, we achieves nearly negligible overhead, as it is fully fused into kernel computation.
@@
$W_{n}$ Coop. Soft Latency (ms) TCs Utilization (%) Valid
1 3.746 10.91
4 0.610 19.71
4 0.613 19.66
BreakDown Analysis. To further analyze the performance gains of BitDecoding, we decompose our optimizations in Table IV. The Lop3 Tensor Cores re-mapping and the asynchronous pipeline contribute to a $1.2\times$$1.5\times$ speedup, while the warp-efficient design achieves more than $2\times$ speedup by enhancing overall computational parallelism.
TABLE IV: Latency (ms) performance with breakdown optimizations.
Optimizations Seq Len
Lop3 Warp Async 32K 64K 128K
0.16 0.31 0.61
0.26 0.50 0.88
0.47 0.85 1.63
0.29 0.53 1.09
0.58 1.08 2.13
@@ VII Related Works
KV cache quantization reduces memory usage in LLMs with long contexts while maintaining performance. Recent works explore 4-bit, 2-bit, and even 1-bit KV cache quantization, aiming to push the limits of compression. Methods like KIVI [17], Gear [12], and KVQuant [11] use per-channel quantization to handle key-value outliers, while RotateKV [25] applies rotation to smooth channel-wise distributions. Although effective at higher compression ratios, these methods lack efficient system implementations, leading to suboptimal performance.
@@
Low-bit weight and low-bit KV cache in LLMs create a unique requirement for mixed-precision matrix multiplication (mpGEMM), where one input matrix is in lower precision (e.g., INT4/2/1) while the other matrix remains in higher precision (e.g., FP16/8). Optimized kernels like Ladder [31] and Marlin [8] improve performance via layout transformations and efficient dequantization. However, these methods require pre-packing and pre-transforming weights, limiting applicability to low-bit KV cache in autoregressive decoding.
@@
KIVI [29] uses Triton with separate kernels for low-bit KV Cache implementation. Atom [35] integrates quantization within the preceding linear layer, while QServe [15] fuses quantization directly into FlashAttention kernels. However, they both rely on GEMV operations with fused multiply–add (FMA) instructions, missing Tensor Core acceleration.
@@ VIII Conclusion
[…]

### v3

@@ IV-A Methods for optimizing low-bit layout on Tensor Cores
(1) Inducing low-bit optimized layout with hardware instructions. Our design is motivated by a novel insight: the thread-to-register mapping of ldmatrix loads data in Tensor Core’s interleaved fragment layout. As shown in Fig. 5-(2), if each thread then quantizes and packs locally, the resulting low-bit packing implicitly preserves the half-precision (FP16) interleaved layout. On unpacking and dequantization, values already match Tensor Core registers—no global reshape is required. Thus, rather than relying on heavyweight global transforms via manual implementations [marlin] or iterative search [ladder] as in prior methods, we use hardware instructions to automatically induce a valid low-bit packing layout while computing. This yields zero-overhead remapping that is efficient, compatible with Tensor Cores execution, and avoids extra data movement.
@@
To mitigate this inefficiency, we further design a faster dequantization mapping approach based on low-level bitwise operations and instructions inspired by [lop3]. After loading packed data into registers using ldmatrix, we cast them to INT32 before mapping them to the interleaved Tensor Core layout following the 75316420 pattern. This layout enables efficient conversion of INT4/INT2 data to FP16 using the lop3 instruction for bitwise manipulation while aligning with the Tensor Core computation pattern.
@@
IV-B Strategies for parallelizing warps
The second challenge is ensuring BitDecoding avoids the pitfalls of existing warp-parallelization strategies for mixed-precision attention, which suffer from low hardware utilization due to frequent warp stalls. Our key insight is that low-bit data moves at much higher bandwidth than full precision, shifting the bottleneck from memory to compute. We therefore design a warp layout that exploits the GPU memory hierarchy to parallelize low-precision operations efficiently, minimizing data movement and substantially improving Tensor Cores utilization (Table III demonstrates minimal overhead).
(1) Enhancing warps parallelism for low-precision operations. We introduce a novel warps layout to enable parallel operations of multiple packed data chunks. Using dequantization as an example, we modify the warp partitioning strategy to better exploit parallelism. As illustrated in Fig. 6, instead of the original strategy that allocates multiple warps along the $M$ dimension, we constrain the allocation to $W_{m}=1$—leveraging the fact that the decoding query length is typically small (<16<16)—and reallocate resources to increase the number of warps along the $N$ dimension ($W_{n}$).
By increasing $W_{n}$, dequantization stalls can be effectively mitigated by the Streaming Multiprocessor (SM) warp scheduler [warp_schedule], as multiple warps concurrently execute dequantization on packed data before proceeding to Tensor Cores-based matrix multiplication.
Similarly, this parallelism strategy alleviates the stalls introduced by on-the-fly quantization in native low-precision attention, ensuring that neither quantization nor dequantization becomes a serialization bottleneck.
@@ V System Implementation
In this section, we describe how we implement BitDecoding, as illustrated in Fig. 7. Our implementation consists of three major components: (i) a query transformation component that supports diverse attention variants in LLMs; (ii) a Residual Kernel that performs low-cost quantization and packing while remaining general to both tensor-wise and channel-wise scaling across quantization algorithms; and (iii) a Packing Kernel with a fine-grained pipeline that fully utilizes both Tensor Cores and CUDA cores. Finally, we discuss architecture-specific optimizations that leverage the advanced features of the latest GPU generations (e.g., Hopper and Blackwell) to further enhance decoding throughput.
@@ V-A Query Transformation
Modern LLMs adopt diverse attention variants [llama3, qwen3, liu2024deepseek] with different key–value (KV) sharing patterns. BitDecoding aims to support all these variants.
@@ V-C Packing Kernel
(1) Optimizing asynchronous data movement. From Global to Shared Memory, we follow FlashAttention [dao2023flashattention2] via block-wise tiling [wang2025tilelang] and strategic recomputation. It processes input matrices $Q\in\mathbb{R}^{T_{m}\times d}$, $K,V\in\mathbb{R}^{T_{n}\times d}$ in tiles within shared memory, using block sizes $T_{m}$ and $T_{n}$. The number of key-value tiles is $C_{n}=\lceil L/T_{n}\rceil$.
@@
From Shared Memory to Register, we use the PTX instruction ldmatrix to efficiently load $K_{\text{pack}}$, $V_{\text{pack}}$ and $sAcc$ from shared memory into registers with the Tensor Cores tiling layout. To eliminate bank conflicts, we use a sizzling scheme [cutlass2024] defined as:
@@
(a) RTX 5090
(b) RTX PRO 6000
Figure 8: Kernel performance with mxfp4 on Blackwell architectures.
V-D Latest Architectures Support
While the design presented thus far effectively targets pre-Hopper architectures (e.g., Ampere), newer generations introduce distinct hardware features that require tailored optimization strategies. Below, we detail how our approach adapts to leverage the specialized instructions and native data formats of the Hopper and Blackwell architectures.
(1) Unlocking Hopper for warpgroup acceleration capabilities via smart uses of PTX-level instructions. Hopper Tensor Cores, increasingly introduce Warpgroup Matrix Multiply-Accumulate (wgmma) instruction. This instruction however imposes a key constraint: in a matrix multiplication $C=AB$, only $A$ and $C$ can be sourced from registers, while $B$ must reside in shared memory. This presents a challenge for low-bit quantized data, as values are typically upconverted to FP16 in registers before computation. To resolve this, we leverage Hopper’s STSM PTX instruction to store dequantized FP16 values in shared memory efficiently, accessible for wgmma_SS operations. Remarkably, the asynchronous nature of WGMMA overlaps storage with computation, optimizing performance.
(2) Accelerating Blackwell with native low-precision format. The Blackwell architecture introduces native support for low-precision tensor operations, eliminating the need for explicit dequantization. Consequently, the lop3-based register remapping described earlier is bypassed in favor of direct execution. We target Blackwell’s low-precision mma instructions—specifically those supporting the micro-scaling formats (e.g., mxfp4 / nvfp4)—to execute GEMM operations directly on packed 4-bit data. While these instructions enforce rigid layout constraints for both the packed values and their block-scaling factors, the layout transformation strategy proposed in Section IV-A is designed to be layout-agnostic. It automatically aligns the packed KV data with the hardware-mandated format, ensuring seamless integration with Blackwell’s native tensor pipelines.
@@ VI Evaluation
In this section, we comprehensively evaluate BitDecoding against state-of-the-art approaches and systems. Our evaluation highlights the following key results:
@@
[…]
@@
    In end-to-end long-context inference, BitDecoding reduces single-batch latency by 3x (on LLaMA-3.1-8B with 128K context) and achieves over 4x higher serving throughput than QServe, demonstrating superior scalability in GQA settings where prior CUDA Core-only methods degrade (Section VI-B).
@@
    BitDecoding preserves near-FP16 accuracy while deriving significant performance gains from each system component, demonstrating only a 0.2% accuracy degradation with 4-bit quantization, while our ablation study confirms that every design module contributes to the overall speedup (Section VI-C).
@@ VI-A Kernels Performance Across GPU Architectures
  • Page: A high-throughput scenario where a larger batch_size is managed using the page management technique [page].
Baselines. We compare BitDecoding against several representative attention kernel implementations. For FP16 KV cache, we use FlashDecoding [dao2023flashattention2, fa3]—a split-partitioned variant of FlashAttention optimized for long-context decoding—as our baseline for speedup normalization. For low-bit KV cache, we evaluate Kivi [kivi], a non-fused kernel supporting 4-bit and 2-bit quantization; Atom [atom] and QServe [qserve], both fused-kernel implementations with CUDA Cores-only approach and supporting 4-bit cache with page management. Notably, Atom does not support GQA.
@@
Figure 9: Kernel performance on Hopper (H100).
[…]
Figure 10: Kernel performance on RTX4090.
[…]
Figure 11: Kernel performance on A100.
Results on Bandwidth-constrained GPU (RTX 4090). Leveraging low-precision data is critical for accelerating inference on bandwidth-constrained GPUs. As shown in Fig. 10, BitDecoding achieves roughly $4\times$ (4-bit) and over $7\times$ (2-bit) speedups over FlashDecoding-v2 in Single and Batches settings, gains that stem directly from alleviating DRAM bottlenecks via low-bit KV caching.
BitDecoding significantly outperforms baselines across all scenarios; unlike the non-fused KIVI, which relies on separate kernels and suffers severe degradation in GQA, BitDecoding’s fully fused design maintains high efficiency. In Page settings, it surpasses fused CUDA-core baselines: for MHA, BitDecoding achieves over $6\times$ speedup compared to QServe’s $3.5\times$. Crucially, in compute-intensive GQA, it maintains a $3\times$ speedup while QServe drops to $1.4\times$, confirming that leveraging Tensor Cores provides robust acceleration where CUDA-only approaches falter.
@@ VI-C Accuracy, Overhead and Performance Breakdown
TABLE I: Efficiency and accuracy tradeoff with low-bit KV cache. We use Llama-3.1-8B-Instruct with seq_len=32Kseq\_len=32K, and evaluate average accuracy on longbench [bai2024longbench].
@@
$W_{n}$ Coop. Soft Latency (ms) TCs Utilization (%) Valid
1 3.746 10.91
4 0.610 19.71
4 0.613 19.66
Figure 14: Runtime overhead of the residual KV cache.
Quantization and Packing Overhead. We evaluate the latency of quantization and packing under a sequence length of seq_len=128Kseq\_len=128K, comparing BitDecoding with Marlin [marlin] and Ladder [ladder]. As shown in Table LABEL:tab:quantization, the pre-transformation and packing step in previous mixed-precision computing methods introduce significant overhead, which cannot be ignored. Our kernel incurs minimal overhead after the Prefill phase, primarily due to kernel launch overhead. Moreover, during decoding, we achieves nearly negligible overhead, as it is fully fused into kernel computation.
@@
BreakDown Analysis. To further analyze the performance gains of BitDecoding, we decompose our optimizations in Fig. 16. Following [ashkboos2024quarot], we use a continuous-packing baseline that quantizes and packs the KV cache at every generation step, which introduces substantial overhead and requires manual effort to maintain valid layouts. In contrast, our layout design automatically induces Tensor Core–compatible layouts for arbitrary low-bit formats, fully unlocking the compute potential of Tensor Cores. On top of this, the warp-parallelism strategy contributes significant additional speedups, while the pipeline optimizations further enhance end-to-end performance.
Figure 16: Breakdown of BitDecoding optimizations across architectural generations.
@@ VII Related Works
KV cache quantization reduces memory usage in LLMs with long contexts while maintaining performance. Recent works explore 4-bit, 2-bit, and even 1-bit KV cache quantization, aiming to push the limits of compression. Methods like KIVI [kivi], Gear [gear], and KVQuant [kvquant] use per-channel quantization to handle key-value outliers, while RotateKV [rotatekv] applies rotation to smooth channel-wise distributions. Although effective at higher compression ratios, these methods lack efficient system implementations, leading to suboptimal performance.
@@
Low-bit weight and low-bit KV cache in LLMs create a unique requirement for mixed-precision matrix multiplication (mpGEMM), where one input matrix is in lower precision (e.g., INT4/2/1) while the other matrix remains in higher precision (e.g., FP16/8). Optimized kernels like Ladder [ladder] and Marlin [marlin] improve performance via layout transformations and efficient dequantization. However, these methods require pre-packing and pre-transforming weights, limiting applicability to low-bit KV cache in autoregressive decoding.
@@
KIVI [triton] uses Triton with separate kernels for low-bit KV Cache implementation. Atom [atom] integrates quantization within the preceding linear layer, while QServe [qserve] fuses quantization directly into FlashAttention kernels. However, they both rely on GEMV operations with fused multiply–add (FMA) instructions, missing Tensor Core acceleration.
@@ VIII Conclusion
BitDecoding establishes a new system foundation for efficient low-bit KV-cache decoding by demonstrating how CUDA cores and Tensor Cores can be cooperatively orchestrated using principled system designs. Its layout-induction and warp-level coordination techniques generalize across attention variants, quantization schemes, and GPU generations, and naturally extend to emerging architectures such as Blackwell and even beyond. We expect BitDecoding to enable future work on algorithm–system co-design for KV-cache quantization, near-lossless test-time scaling, and more capable GPU execution models for long-context LLMs inference.